# Exact Expressive Power of Transformers with Padding

### v1

@@ 5 Logd-Looped Padded Transformers Recognize Exactly $\mathsf{TC}^{d}$
Lemma 4 directly provides one direction. To show the other direction that $\mathsf{L}$-uniform $\mathsf{TC}^{d}\subseteq\mathsf{AHAT}^{d}_{*}$, we will apply Lemma 3 with the wide-$\mathsf{TC}^{d}$ circuit evaluation problem (Appendix B) as $L$, $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ as class $\mathsf{C}$, and $\mathsf{L}$ as class $\mathsf{R}$. We next argue that the preconditions of Lemma 3 are met, finishing the proof. First, 7.1 (Appendix B) shows that logd-depth looped transformers (without padding) can solve the wide-$\mathsf{TC}^{d}$ circuit evaluation problem, i.e., $L\in\mathsf{AHAT}^{d}_{0}$. Combining this with Lemma 4, we also obtain that $L\in\mathsf{L}\textrm{-uniform}\ \mathsf{TC}^{d}$. Lemma 6 (Appendix B) shows that $L$ is hard for $\mathsf{L}$-uniform $\mathsf{TC}^{d}$ under $\mathsf{L}$ reductions. Together, we obtain that $L$ is complete for this class. Finally, Theorem 2 implies that logd-depth transformers can recognize any language in $\mathsf{L}$, and thus compute any $\mathsf{L}$ reduction in the sense of Definition 10. Applying Lemma 3, we conclude that $\mathsf{L}$-uniform $\mathsf{TC}^{d}\subseteq\mathsf{AHAT}^{d}_{*}$. ∎
6 Conclusion
Our results in this work give a precise theoretical understanding of how padding and looping—two ways to dynamically expand the computational resources of a transformer at inference time—increase the expressive power of transformers. Padding expands the circuit width of transformers, allowing them to resolve logical formulas over more variables. As a consequence of this, polynomially padded transformers can recognize exactly $\mathsf{TC}^{0}$, which was previously known only as an upper bound (Merrill and Sabharwal, 2023a; Pfau et al., 2024). In contrast, looping increases the depth of transformers. Applying looping on top of padding, we extended our result to show that logd-depth, padded transformers recognize exactly $\mathsf{TC}^{d}$. This means that transformers with polynomial padding and polylogarithmic looping converge to recognizing $\mathsf{NC}$, the largest class of problems that can be solved with parallel computation. In contrast, transformers with CoT have greater expressive power under standard complexity conjectures (Merrill and Sabharwal, 2023b), but suffer from slow sequential decoding. Thus, while looping and padding are not as powerful as CoT, our results suggest they are quite effective ways to expand transformers’ expressive power while preserving parallelism.
Several interesting open questions remain from this work. On the theoretical side, it would be valuable to develop a more finegrained characterization of $\mathsf{AHAT}^{d}_{k}$, i.e., looped transformers where the padding is at most $O(n^{k})$ for some fixed $k$. On the empirical side, while looped and padded transformers have both been explored already to an extent, it would be interesting to see whether these approaches could be successfully integrated to improve the performance of transformers on hard reasoning tasks. Further developing these approaches could ultimately lead to inference-time compute methods based on padding and looping that increase transformers’ expressive power without sacrificing parallelism, providing a more efficient alternative to CoT for solving moderately parallelizable problems.
Acknowledgments
Thanks to Selim Jerad, Andy Yang, and Michael Cadhilac, and attendees at the Formal Languages and Neural Networks (FLaNN) seminar for discussions. This project was supported by the National Science Foundation (NSF) through award 1922658 and WM’s NSF Graduate Research Fellowship. WM was also supported by a Two Sigma PhD Fellowship as well as by the Allen Institute for AI.
Appendix A Omitted Proofs
@@
That unmasked padded transformers can be simulated by mixed-masked transformers of the same depth holds trivially, as the former constitute a special case of the latter. We will next describe how to convert an unmasked padded transformer to a masked one, on a per-head basis. The construction will leave the computation of masked heads unchanged, making the approach suitable for converting both unmasked and mixed-masked padded transformers to masked ones.
Let $E$ be an unmasked encoder in $\mathsf{uAHAT}^{d}_{k}$. Then $E$ has depth $\ell=O(\log^{d}n)$ and operates over $n+O(n^{k})$ tokens (including original input and padding tokens). By Lemma 1, $E$ can be simulated by a causally masked decoder $D$ with depth $\ell$ and with $\ell\cdot(n+O(n^{k}))$ new padding tokens (after the original $O(n^{k})$ padding tokens used by $E$). Thus, the total number of padding tokens $D$ uses is $p=O(n^{k})+\ell\cdot(n+O(n^{k}))$. Since $\ell=O(\log^{d}n)$, this simplifies to $p=O(\log^{d}(n)\cdot n^{k^{\prime}})$ where $k^{\prime}=\max\{k,1\}$. This is $O(n^{k^{\prime}})$ when $d=0$, and $O(n^{k^{\prime}+1})$ when $d\geq 1$. This finishes the proof of the first two parts. For the third part, observe that by definition, $\mathsf{uAHAT}^{d}_{*}=\bigcup_{k=0}^{\infty}\mathsf{uAHAT}^{d}_{k}$. From the above, this in turn is contained in $\bigcup_{k=0}^{\infty}\mathsf{AHAT}^{d}_{k^{\prime}+1}\subseteq\mathsf{AHAT}^{d}_{*}$, as desired.
For the mixed-masking case, the construction in Lemma 1 will preserve the original computation of causally masked heads if we add an additional term with large negative weight $C$ (fixed w.r.t. $n$) that is activated if the index within the block is greater than $b_{i}$. We set $C$ large enough to dominate all other terms in the inner product computation at each token. As a result, the head is constrained to only attend to previous tokens within the block. Applying this modified construction on a case-by-case basis per head, we can take Proposition 1 to apply equally to transformers with mixed-masking. ∎
See 4
@@
As noted earlier, it is known that $\mathsf{AHAT}^{0}\subseteq\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{0}$ [Merrill and Sabharwal, 2023a, c, Chiang, 2025], and this generalizes to transformers with padding tokens [Pfau et al., 2024]. We will extend this construction to the case of $\mathsf{AHAT}^{d}_{*}$. The main observation is that even though the depth of a $O(\log^{d}n)$-depth padded AHAT transformer is input-dependent, the transformer has a looped architecture and thus a fixed (independent of input size) set of parameters. This makes it possible for a Turing machine (which can only store a fixed number of parameters) to construct an equivalent $\mathsf{TC}^{d}$ circuit family. [^3] The proof naturally generalizes to the case of logd-depth transformers that don’t necessarily have identical parameters across blocks of layers but whose parameters in each layer can be computed in an $\mathsf{L}$-uniform manner.
@@
Let $L$ be a language in $\mathsf{AHAT}^{0}_{*}$ and let $T$ be a looped, padded AHAT transformer that, when unrolled $c\log^{d}n$ times for large enough $n$, [^4] As before, for small $n$, we assume $T$ uses a lookup table. recognizes whether $x\in\mathsf{L}$ for any input $x$ with $\lvert x\rvert=n$. Let $\langle A,B,C\rangle$ be the partition of layers of $T$ where $A$ is the set of initial layers, $B$ is the block that’s repeated $c\log^{d}n$ times on inputs of length $n$, and $C$ is the set of final layers. Each of these itself is a fixed-depth padded transformer; let’s call these $T_{A}$, $T_{B}$, and $T_{C}$. By prior results, there are $\mathsf{L}$-uniform $\mathsf{TC}^{0}$ circuit families $\{C^{A}_{n}\}_{n=0}^{\infty},\{C^{B}_{n}\}_{n=0}^{\infty},$ and $\{C^{B}_{n}\}_{n=0}^{\infty}$ that simulate transformers $T_{A},T_{B},$ and $T_{C}$, respectively. Thus, there is a (single) log-space Turing machine $M$ that, on input $\lvert x\rvert$, can construct each of these three circuit families. By looping the $T_{B}$ block $c\log^{d}n$ times, it can therefore also construct a circuit family $\{C_{n}\}_{n=0}^{\infty}$ that recognizes $L$. It follows that $L$ is in $\mathsf{L}$-uniform $\mathsf{TC}^{d}$. ∎
This argument does not work for $\mathsf{FO}$-uniform $\mathsf{TC}^{d}$ since it is not clearly that case that the circuit that repeats a layer $\log n$ times can be defined in $\mathsf{FO}$ just because the one that simulates one layer can [Merrill and Sabharwal, 2023c].
Appendix B $\mathsf{TC}^{d}$ Circuit Evaluation Problem
[…]
$\mathrm{Circuit}$ $\to\ \mathrm{Gate}^{*}$
$\mathrm{Gate}$ $\to\ \texttt{X}\;\mid\;\mathrm{Op}\;\mathrm{Arg}^{*}$
[…]
$\mathrm{Arg}$$\to\ \texttt{\&}\texttt{1}^{*}$

### v2

@@ Appendix B Uniformity of Circuit Classes
To use the connection language notion of uniformity to serialize the circuit with log space, we simply maintain a counter of our current gate position and edge and call an $\mathsf{L}$ oracle for the connection language to print each gate and wire. In the other direction, we modify the routine that serializes the circuit to only output gate $i$ or edge $i,j$, and we use this to recognize $L_{\mathcal{C}}$. ∎
B.1 Uniformity Composition Closures
Our later results about uniform circuit families will rely on the property (formalized below) that for strong-enough classes $\mathsf{A}$, any fixed composition of $\mathsf{A}\textrm{-uniform}$ circuit families is also $\mathsf{A}\textrm{-uniform}$. Note that any fixed composition of functions can be decomposed into a fixed number of serial and parallel compositions of two functions, defined as follows:
Definition 14 (Function Composition).
For functions $f,g\in\{0,1\}^{*}\to\{0,1\}^{*}$, their serial composition is the function $(g\circ f)(w)=g(f(w))$ and their parallel composition is the function $\langle f,g\rangle(w)=f(w)\cdot g(w)$.
For $\mathsf{L}\textrm{-uniform}$ functions, serial and parallel composition clearly preserves $\mathsf{L}\textrm{-uniformity}$, as we can maintain a counter to reroute inputs and outputs between subcircuits appropriately. The same applies to any class $\mathsf{A}$ containing $\mathsf{L}$. For $\mathsf{FO}$ uniform circuit classes, closure under composition requires a bit more work to justify, as shown in the following two lemmas.
Lemma 7.
If functions $f$ and $g$ have polynomial-size $\mathsf{FO}\textrm{-uniform}$ circuit families, then so does their parallel composition $\langle f,g\rangle$.
Proof.
Let $\mathcal{C}^{f}=\{C^{f}_{n}\}_{n=0}^{\infty}$ be an $\mathsf{FO}\textrm{-uniform}$ circuit family for $f$, with connection language $L_{f}=L_{\mathcal{C}_{f}}\in\mathsf{FO}$. Similarly, define $\mathcal{C}^{g},C^{g}_{n},$ and $L_{g}$ for $g$.
We will decide the connection language $L_{\langle f,g\rangle}$ by querying $L_{f}$ and $L_{g}$ to build $C^{f}_{n}$ and $C^{g}_{n}$ in parallel. Without loss of generality, we assume all gate indices in the input word for a gate query fit into one of the following three cases: (1) at most the number of gates $s_{f}$ in the circuit $C^{f}_{n}$, (2) larger than $s_{f}$ but at most $s_{f}+s_{g}$ where $s_{g}$ is the number of gates in $C^{g}_{n}$, or (3) larger than $s_{f}+s_{g}$ but at most $s_{f}+s_{g}+o_{f}$, where $o_{f}$ is the number of output gates of $C^{f}_{n}$. If all gates are at most $s_{f}$, we simply query $L_{f}$. If all gates are larger than $s_{f}$ but at most $s_{f}+s_{g}$, we subtract $s_{f}$ from each gate index and query $L_{g}$. Finally, if we are querying new gates from $s_{f}+s_{g}<i\leq s_{f}+s_{g}+o_{f}$, we design the connection language to represent identity gates. Importantly, addition and inequality checks can be performed in $\mathsf{FO}$, and the connection languages $L_{f}$ and $L_{g}$ can be queried in $\mathsf{FO}$ by construction.
The logic for edge queries in the connection language modifies gate indices according to the same logic as the gate queries with a few exceptions. First, for an edge $(i,j)$, if $i$ is an input for $C^{f}_{n}$ (and $C^{g}_{n}$) and $i=j-s_{f}$, we return $1$. This ensures that $C^{f}_{n}$ and $C^{g}_{n}$ both receive the same input. Second, if $i$ is an output for $C^{f}_{n}$ and $s_{f}+s_{g}<j\leq s_{f}+s_{g}+o_{f}$, we also return 1. This copies over the outputs of $C^{f}_{n}$ so that they appear at the very end of the combined circuit.
Thus, this connection language constructs a circuit family that computes $f$ and $g$ in parallel and returns their outputs $\langle g(w),f(w)\rangle$. Without loss of generality, their order can be easily permuted to obtain the output $\langle f(w),g(w)\rangle$ of parallel composition. It is clear from the construction that the size of the resulting circuit family is linear in the sizes of $\mathcal{C}^{f}$ and $\mathcal{C}^{g}$, and thus polynomial in $n$. ∎
Lemma 8.
If functions $f$ and $g$ have polynomial-size $\mathsf{FO}\textrm{-uniform}$ circuit families, then so does their serial composition $g\circ f$.
Proof.
Let $\mathcal{C}^{f}=\{C^{f}_{n}\}_{n=0}^{\infty}$ be an $\mathsf{FO}\textrm{-uniform}$ circuit family for $f$, with connection language $L_{f}=L_{\mathcal{C}_{f}}\in\mathsf{FO}$. Similarly, define $\mathcal{C}^{g},C^{g}_{n},$ and $L_{g}$ for $g$.
We will decide the connection language $L_{g\circ f}$ by querying $L_{f}$ to construct $C^{f}_{n}$, then modifying the inputs to $L_{g}$ to build $C^{g}_{m}$ on the output of $f$, which has size $m$. The cases are the same as for parallel composition, except for a few changes. First, we build $C^{g}_{m}$ instead of $C^{g}_{n}$. Second, when querying input gates for $C^{g}_{m}$, we add a condition that routes from the outputs of $C^{f}_{n}$ rather than its inputs. Finally, we do not construct additional gates to copy over the outputs from $C^{f}_{n}$. Thus, this connection language builds a circuit family that computes $g$ on the output of $f$. Similar to the parallel composition case, it is clear from the construction that the size of the resulting circuit family is linear in the sizes of $\mathcal{C}^{f}$ and $\mathcal{C}^{g}$, and thus polynomial in $n$. ∎
Combining Lemmas˜7 and 8 along with the fact that any fixed composition of functions can be decomposed into a fixed number of serial or parallel compositions of two functions at a time, we obtain that fixed function composition preserves $\mathsf{FO}\textrm{-uniformity}$ of circuit families:
Proposition 3 (Composition Preserves $\mathsf{FO}$-Uniformity).
Any fixed composition of polynomial-size $\mathsf{FO}\textrm{-uniform}$ circuit families is also polynomial-size and $\mathsf{FO}\textrm{-uniform}$.
Lemma 9 (Recurrent Composition).
Let $m(n)$, $d(n)$, and $r(n)$ be functions at most polynomial in $n$, with $m(n)$ and $r(n)$ definable as variables in $\mathsf{FO}$ given $a^{n}$. Let $f:\{0,1\}^{m(n)}\to\{0,1\}^{m(n)}$ have a polynomial-size $\mathsf{FO}$-uniform circuit family with depth $d(n)$. Then the function $f^{r(n)}$ (i.e., $f$ called recurrently on itself $r(n)$ times) has an $\mathsf{FO}$-uniform circuit family of polynomial size and depth d(n)r(n)d(n)\;r(n).
Proof.
Let $\mathcal{C}^{f}=\{C^{f}_{n}\}_{n=0}^{\infty}$ be a polynomial size $\mathsf{FO}$-uniform circuit family for $f$, with connection language $L_{f}$. Without loss of generality, we assume the number of gates in $C^{f}_{n}$ is $2^{s(n)}$ for some integer $s(n)$ (this can always be achieved by padding $C^{f}_{n}$ if necessary, in a way that increases size by at most a factor of 2). Since the circuit family is of polynomial size, $s(n)=\mathrm{O}(\log n)$ and can be computed in $\mathsf{FO}$ as the smallest $j$ s.t. $\mathsf{bit}(s_{f},j)=1$, where $s_{f}$ is the number of gates in $C^{f}_{n}$. We will construct an $\mathsf{FO}$-uniform circuit family $\mathcal{C}=\{C_{n}\}_{n=0}^{\infty}$ where $C_{n}$ consists of $f(n)$ iterations of $C^{f}_{n}$. By construction, $C_{n}$ has size at most polynomial and depth d(n)r(n)d(n)\;r(n).
The remainder of the proof will justify that $\mathcal{C}$ is $\mathsf{FO}$-uniform by defining its connection language $L_{\mathcal{C}}$ in terms of $L_{f}$, starting with gate queries. Given a gate query $w$ with index $i$, we first compute $r(n)$, which is possible by construction. We then compute $S=r(n)\cdot 2^{s(n)}$ by left shifting $r(n)$ by $s(n)$, and, if $i\geq S$, we reject $w$. Otherwise, we compute $i^{\prime}=i\mod 2^{s(n)}=\mathsf{bit}(i,s(n))$, which can be computed in $\mathsf{FO}$ by reading the $s(n)$-th bit from the start of $i$ in $w$. We then compute $u$ as a new query where $i$ is replaced by $i^{\prime}$. We then query whether $u\in L_{f}$. This ensures that $C_{n}$ repeats all the gates in $C^{f}_{n}$ stacked in $r(n)$ blocks.
Given an edge query $w$ between $(i,j)$, we first compute $i^{\prime}=i\bmod 2^{s(n)}$ and $j^{\prime}=j\mod 2^{s(n)}$ as above. Additionally, let $q_{i}=\lfloor i/2^{s(n)}\rfloor$ and $q_{j}=\lfloor j/2^{s(n)}\rfloor$. We compute $q_{i}$ and $q_{j}$ in $\mathsf{FO}$ by right-shifting by $s(n)$. If $q_{i}=q_{j}$, we follow similar logic to the gate case, constructing $u$ by replacing $i$ with $i^{\prime}$ and $j$ with $j^{\prime}$. We then query whether $u\in L_{f}$. This has the effect of constructing all edges within a block of $C_{n}$ analogously to those in $C^{f}_{n}$. Additionally, if $q_{i}+1=q_{j}$, $i^{\prime}\geq s_{f}-m(n)$, and $j^{\prime}\leq m(n)$, we return a 1 for this edge; recall that $m(n)$ is the number of inputs as well as the number of outputs of $f$. This has the effect of routing the output of each block as the input for the next block. Thus, $\mathcal{C}$ computes $f^{r(n)}$ and is $\mathsf{FO}$-uniform. ∎
Appendix C Omitted Proofs
@@
As noted in the proof of Lemma˜1, the $1/i$ position encoding used in our unmasked transformers can be computed by causally masked transformers, and hence also by mixed-masked transformers. Thus, unmasked padded looped transformers with position encoding $1/i$ constitute a special case of mixed-masked padded looped transformers, and can thus be trivially simulated by the latter. We will next describe how to leverage Lemma˜1 to convert an unmasked padded looped transformer to a masked one, on a per-head basis. The construction will leave the computation of masked heads unchanged, making the approach suitable for converting both unmasked and mixed-masked padded transformers to masked ones.
Let $E$ be an unmasked looped encoder in $\mathsf{uAHAT}^{d}_{k}$. Then $E$ has depth $\ell=O(\log^{d}n)$ and operates over $n+O(n^{k})$ tokens (including original input and padding tokens). By Lemma˜1, $E$ can be simulated by a causally masked decoder $D$ with depth $\ell$ and with $\ell\cdot(n+O(n^{k}))$ new padding tokens (after the original $\mathrm{O}(n^{k})$ padding tokens used by $E$). Thus, the total number of padding tokens $D$ uses is $p=O(n^{k})+\ell\cdot(n+O(n^{k}))$. Since $\ell=O(\log^{d}n)$, this simplifies to $p=O(\log^{d}(n)\cdot n^{k^{\prime}})$ where $k^{\prime}=\max\{k,1\}$. This is $\mathrm{O}(n^{k^{\prime}})$ when $d=0$, and $\mathrm{O}(n^{k^{\prime}+1})$ when $d\geq 1$. This finishes the proof of the first two parts. For the third part, observe that by definition, $\mathsf{uAHAT}^{d}_{*}=\bigcup_{k=0}^{\infty}\mathsf{uAHAT}^{d}_{k}$. From the above, this in turn is contained in $\bigcup_{k=0}^{\infty}\mathsf{AHAT}^{d}_{k^{\prime}+1}\subseteq\mathsf{AHAT}^{d}_{*}$, as desired.
For the mixed-masking case, the construction in Lemma˜1 will preserve the original computation of causally masked heads if we add an additional term with large negative weight $C$ (fixed w.r.t. $n$) that is activated if the index within the block is greater than $b_{i}$. We set $C$ large enough to dominate all other terms in the inner product computation at each token. As a result, the head is constrained to only attend to previous tokens within the block. Applying this modified construction on a case-by-case basis per head, we can take this proposition to apply equally to transformers with mixed-masking. ∎
The above simulation of unmasked padded looped transformers (with position encoding $1/i$) with mixed-mask transformers can, in fact, be extended even to the case where the unmasked transformer uses $i/n$ positional encoding:
Lemma 10.
There exists a mixed-mask sublayer with one masked head and one unmasked head that computes $\phi(i/n)$.
@@
We use two attention heads, one causally masked and one unmasked, both of which attend uniformly with value $1$ only for the beginning-of-sequence symbol. The first head thus computes $1/i$, while the second head computes $1/n$. We can then combine these values to compute $\phi(1/n,1/i)=\phi(i/n)$. ∎
See 3
Proof.
Consider any language $L^{\prime}\in\mathsf{C}$. By the assumed completeness of $L$, there exists an $\mathsf{R}$ reduction $f$ that maps inputs of $L^{\prime}$ into inputs of $L$ such that $w\in L^{\prime}$ if and only if $f(w)\in L$. From the last precondition of the theorem, there exists a causally masked logd-depth padded transformer $T_{f}$ that recognizes the corresponding reduction language $R_{f}$ or, equivalently, computes the reduction function $r_{f}$ (cf. Definition˜10). We will assume the latter, i.e., that $T_{f}$ computes $r_{f}$, though the construction can also be made to work if $T_{f}$ checks membership in $R_{f}$. Additionally, we also have from a precondition that there exists a causally masked logd-depth padded transformer $T_{L}$ that recognizes $L$.
Compute Reduction. Suppose $T_{f}\in\mathsf{AHAT}^{d}_{k}$ and let $n=\lvert w\rvert$. Then, on input $(w,i)$, where $i$ is represented in binary, $T_{f}$ uses $\mathrm{O}(n^{k})$ padding tokens to compute $f(w)_{i}$, which we upper bound by $n^{k+1}$ for sufficiently large $n$ (for small $n$, we assume $T_{f}$ instead uses a fixed lookup table (and no padding). There is a uniform $\mathsf{TC}^{0}$ circuit family that computes $\mathsf{bit}(i,j)$ from input $\phi(i),\phi(j)$, so, by Theorem˜1, there exists a transformer $T_{\mathsf{bit}}$ that computes $\mathsf{bit}$ with $n^{b}$ padding tokens, for some $b$ and sufficiently large $n$ (again, for smaller $n$, we assume $T_{\mathsf{bit}}$ uses a fixed lookup table). Since $\lvert f(w)\rvert$ is bounded by some polynomial $n^{c}$, we know that any $i\leq\lvert f(w)\rvert$ takes at most $c\log n$ bits to specify, which we upper bound by $c^{\prime}=n$ (since it’s unclear how a transformer would compute $\log n$).
We will construct $n^{c}$ blocks of padding tokens, each of size $B=c^{\prime}n^{b}+n^{k+1}$. Block $i$ will first compute the binary expansion of $i$ in its first $c^{\prime}n^{b}$ padding tokens using $T_{\mathsf{bit}}$, and then use the remaining $n^{k+1}$ padding tokens to consume $(w,i)$ using $T_{f}$ and return $r_{f}(w,i)$. At each token $t$ in block $i$, we compute $\phi(i)$ as $\phi(\lfloor t/B\rfloor)$, using the fact that we can compute integer division with a fixed block of transformer layers [merrill-2025-little] and having computed $B$ as a function of $n$ in earlier layers (since it is in $\mathsf{TC}^{0}$). For $t^{\prime}\leq c^{\prime}$, token $t=t^{\prime}n^{b}$ computes $\phi(t^{\prime})$ via division similarly to $\phi(i)$ and then $\mathsf{bit}(i,t^{\prime})$. This recovers the binary representation of $i$ stored across tokens $t=t^{\prime}n^{b}$ in block $i$, for $1\leq t^{\prime}\leq c^{\prime}$. Finally, we apply a slightly modified $T_{f}$ over the block. Specifically, we add a new term to each attention head so that it only attends over the input tokens and some tokens within the current block: those satisfying $t=t^{\prime}n^{b}$ for some $1\leq t^{\prime}\leq c^{\prime}$ or $t>c^{\prime}n^{b}$: since $c^{\prime}=n$, these predicates are simple to check. As a result, block $i$ simulates $T_{f}$ over $w$ with $n^{k}$ padding. Thus, the final token of block $i$ computes $r_{f}(w,i)=f(w)_{i}$.
See 6
Proof.
@@
Let $L$ be a language in $\mathsf{AHAT}^{d}_{*}$ and let $T$ be a looped, padded AHAT transformer that, when unrolled $c\log^{d}n$ times for large enough nn, [^3] As before, for small $n$, we assume $T$ uses a lookup table. recognizes whether $w\in L$ for any input $w$ with $\lvert w\rvert=n$. Let $\langle A,B,C\rangle$ be the partition of layers of $T$ where $A$ is the set of initial layers, $B$ is the block that’s repeated $c\log^{d}n$ times on inputs of length $n$, and $C$ is the set of final layers. Each of these itself is a fixed-depth padded transformer; let’s call these $T_{A}$, $T_{B}$, and $T_{C}$. By prior results [merrill-sabharwal-2023-parallelism, merrill2023logic, chiang2025transformers], there are $\mathsf{FO}\textrm{-uniform}$ $\mathsf{TC}^{0}$ circuit families $\{C^{A}_{n}\}_{n=0}^{\infty},\{C^{B}_{n}\}_{n=0}^{\infty},$ and $\{C^{B}_{n}\}_{n=0}^{\infty}$ that simulate transformers $T_{A},T_{B},$ and $T_{C}$, respectively. Let $T_{R}$ be $T_{B}$ iterated $\log n$ times, and Let $T^{d}_{R}$ be $T_{B}$ iterated $\log^{d}n$ times. We justify that $r(n)=\lceil\log n\rceil$ is definable as a variable in $\mathsf{FO}$ given $a^{n}$: compute the index of the last $a$ and then find the greatest bit index that is 1. We now invoke Lemma˜9 to show that $T_{R}$ has an $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ circuit family; we can repeat this process $d$ times to get an $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ circuit family for $T^{d}_{R}$. By Lemma˜8, there is a also an $\mathsf{FO}\textrm{-uniform}\ \mathsf{TC}^{d}$ circuit family that computes the serial composition of $T_{A},T_{R},\ldots,T_{R},$ and $T_{C}$, where $\ldots$ accounts for a fixed repetition of $T_{R}$ to account for a constant $c$ on the depth $c\lceil\log n\rceil^{d}$. Thus, we conclude that $L\in\mathsf{FO}\textrm{-uniform}\mathsf{TC}^{d}$. ∎
Appendix D Wide-$\mathsf{TC}^{d}$ Circuit Evaluation Problem
To formalize the circuit evaluation problem, we will use the following serialized format for representing a circuit. This format is a simplification of the one used by merrill-sabharwal-2023-parallelism, with two main differences: (a) instead of a single threshold gate, we use $\texttt{AND},\texttt{OR},\texttt{NOT},$ and MAJ (the majority gate), which will simplify the description of the construction; and (b) we do not require the gates in the serialization to be sorted in any particular order, which makes it easier (and perhaps even possible) to have a log-space reduction from any $\mathsf{L}\textrm{-uniform}$ $\mathsf{TC}^{d}$ language to the circuit evaluation problem in this specific serialization format. The syntax of this circuit format is governed by the following grammar:
$\mathrm{Circuit}\$$\to\ \mathrm{Gate}^{*}$
$\mathrm{Gate}\$$\to\ \texttt{X}\;\mid\;\mathrm{Op}\;\mathrm{Arg}^{*}$
$\mathrm{Op}\$$\to\ \texttt{AND}\;\mid\;\texttt{OR}\;\mid\;\texttt{NOT}\;\mid\;\texttt{MAJ}$
$\mathrm{Arg}\$$\to\ \texttt{\&}\texttt{1}^{*}$